# Semiparametric Discovery and Estimation of Interaction in Mixed Exposures using Stochastic Interventions

### v2

@@ 5.1 Evaluating Performance
We assessed the asymptotic convergence to the true exposure relationships used in the DGP, as well as the convergence to the true counterfactual differences for these exposures, in each simulation. To do so, we followed the following steps 1. From this sample, we generated a random sample of size $n$ from the DGP, 2. At each iteration, we used the training sample to define the exposure(s) and create the necessary estimators for the target parameter dependent on the variable sets. We then used the validation sample to obtain the updated causal parameter estimate using TMLE. We repeated this process for all the folds, 3. At each iteration, we output the stochastic shift estimates given the pooled TMLE.
To evaluate the performance of our approach, we calculated several metrics for each iteration, including bias, variance, MSE, confidence interval (CI) coverage, and the proportion of instances in which the true variable relationships were identified. To visually inspect if the rate of convergence was at least as fast as $\sqrt{n}$, we show projections of a $\sqrt{n}$ consistent estimator starting from the initial bias. We then calculated the variance for each estimate and used it to calculate the mean squared error (MSE) as $\text{MSE}=\text{bias}^{2}+\text{variance}$. To account for different sample sizes, we multiply the MSE estimates by $n$. For each estimate, we also calculated the CI coverage of the true stochastic shift parameter given the exposure determined adaptively from the data. We calculated these performance metrics at each iteration, performing 50 iterations for each sample size $n=$ (250, 500, 1000, 1500, 2000, 2500, 3000, 5000). We used SuperNOVA with 5-fold cross-validation (to speed up calculations in the simulations) and default learner stacks (see below) for each nuisance parameter and data-adaptive parameter. Additionally, the quantile threshold to filter the basis functions based on F-statistic was set to 0 to include all basis functions used in the final best-fitting model. For each sample size, we calculate the proportion of iterations in which the correct main effects and interaction effects were discovered.
To use SuperNOVA, we need estimators for $\bar{Q}=E(Y|A,W)$ and $g_{n}=p(A|W)$ (conditional density). SuperNOVA provides default algorithms to be used in a Super Learner van der Laan Mark et al. [2007] that are both fast and flexible. For our data-adaptive procedure, we include learners from the packages earth Milborrow. Derived from mda:mars by T. Hastie and R. Tibshirani. [2011], polspline Ripley and Venables [2021], and hal9001 Coyle et al. [2022]. The results from each of these packages can be formed into a model matrix, on which we can fit an ANOVA to obtain the resulting linear model of basis functions. To estimate $\bar{Q}$, we include estimators for the Super Learner from glm, elastic net Tay et al. [2023], random forest Wright and Ziegler [2017], and xgboost Chen and Guestrin [2016]. For the semiparametric density estimator $g_{n}$, we create estimators based on homoscedastic errors (HOSE) and heteroscedastic errors (HESE) from the same estimators used in $\bar{Q}$.
@@ 5.2 Results
SuperNOVA consistently discovers interactions. The interaction effect was identified in 100% of the folds in all sample sizes. This indicates that, at least in this simple simulation, the default learners are flexible enough to capture the interaction effect. Target parameters estimated by SuperNOVA converge to truth at a $1/\sqrt{n}$ rate, Figure 1 A shows the absolute bias, CI coverage, mean squared error, and standard deviation as the sample size increases to 5000. For bias, MSE and standard deviation, there is a converge to zero at sample size 5000. For coverage, the average coverage for each target parameter was: individual shift: 98%; joint shift: 96%; interaction: 96%.
Figure 1: A. Bias, CI, Standard Deviation and MSE for Each Parameter Across Sample Sizes: These metrics represent results from CV-TMLE, that is bias is w.r.t. the data adaptive target parameter where we use CV to identify the exposure and then estimate. B. Bias Standardized by Standard Error Compared to Ground-Truth Outcome Under Shift Interventions
To establish SuperNOVA’s estimator validity, it is essential to demonstrate that its sampling distribution approximates a normal distribution centered at 0, with its variance decreasing as sample size grows. This is evaluated through the sampling distribution of each estimate. Figure 1 B illustrates the probability density of the standardized bias relative to true estimates in 50 iterations for varying sample sizes, indicating convergence to a mean 0 normal distribution as sample size increases. Specifically, Figure 1 B. Individual Effect displays the convergence of the standardized bias for a single variable shift towards a normal distribution with mean 0 and standard deviation 1. Likewise, Figure 1 B. Joint Effect and Figure 1 B. Interaction depict the sampling distributions for dual variable shifts and interaction parameter estimates, respectively. In all scenarios, the distributions align with a standard normal distribution as the sample size increases, ensuring the reliability of our inferential approach.
@@ 6.1 Analysis of NIEHS Synthetic Mixtures Data
The NIEHS synthetic mixtures data is a commonly used data set to evaluate the performance of statistical methods for mixtures. These synthetic data can be considered the results of a prospective cohort study, where the outcome cannot cause exposures, and correlations between exposure variables can be thought of as caused by common sources or modes of exposure. The variable $W$ can be assumed to be a potential confounder and not a collider. The data set has 7 exposures ($A_{1}-A_{7}$) with a complex dependency structure based on endocrine disruption. Two groups of exposure ($A_{1},A_{2},A_{3}$ and $A_{5},A_{6}$) lead to high correlations within each group. $A_{1},A_{2},A_{7}$ positively contribute to the outcome, $A_{4},A_{5}$ have negative contributions, while $A_{3}$ and $A_{6}$ have no impact on the outcome. Rejecting $A_{3}$ and $A_{6}$ is difficult due to their correlations with the members of the cluster group. This correlation and effects structure is biologically plausible, as different congeners of a group of compounds may be highly correlated but have different biological effects. Exposures have various agonistic and antagonistic interactions, a breakdown of the variables sets and their relationships are: $A_{1},A_{2}$ both increase Y by concetration addition, $A_{1}-A_{4},A_{2}-A_{4},A_{1}-A_{4},A_{2}-A_{4}$ competetive antagonism, $A_{1}-A_{7},A_{2}-A_{7}$ supra-additive or synergistic, $A_{4},A_{5}$ both addivively decrease Y, $A_{4}-A_{7},A_{5}-A_{7}$ unusual antagonism. Synthetic data and the key for dataset 1 are available on GitHub. This resource shows the interactions and marginal dose relationships built into the data. Given these toxicological interactions, we expect these sets of variables to be determined in SuperNOVA. For example, we might expect a positive counterfactual result for A1, A2, A7 and negative results for A4 and A5. Likewise, in the case for antagonistic relationships such as A1-A5 (third row in Table 1), we would expect a joint shift to come closer to the null, since A5 antagonizes the positive effects of A1. For A1 and A2, we would expect the joint shift to be close to the sum of individual shifts (not much interaction), but for A1 and A7 we expect to estimate more than an additive effect (some interaction). The NIEHS data set has 500 observations and 9 variables. $W$ is a binary confounder. Of course, in this data there is no ground truth, like in the above simulations, but we can gauge SuperNOVA’s performance by determining if the correct variable sets are used in the interactions and if the correct variables are rejected.
We apply SuperNOVA to this NIEHS synthetic data using a 20-fold CV. For computational simplicity, we considered a small learner library consisting of a main terms linear regression and random forest models for estimation $g$ and $Q$. For the discovery of the exposure variable sets, we built 13 earth models with various hyperparameters. We parallelize over the cross-validation to test computational run-time on a newer personal machine an analyst might be using. We use a delta of 1 for all the exopsures in the mixture. We present the pooled TMLE results and the pooled TMLE results with inverse variance for null folds. Due to space limitations we do not show fold specific results but these are provided to the user.
@@ 6.2 NIEHS Synthetic Data Results
SuperNOVA accurately rejects exposures $A_{3},A_{6}$, finding the correct marginal effects for $A_{1},A_{2},A_{4},A_{5},A_{7}$ consistently in all folds. Specifically, A1: 100%, A2: 85%, A4: 95%, A5: 100%, A7 100%. The interactions were harder to identify consistently, and the strongest interaction between A5 and A7 was detected in 70% of the folds. Specifically, A1-A2: 10%, A1-A5: 45%, A1-A7: 20%, A2-A7: 50%, A4-A5: 15%, A4-A7: 5%, and A5-A7: 70%. Our analysis of the NIEHS Mixtures Data reveals the marginal effects of various exposures on endocrine disruption outcomes, measured as the change in outcome per unit increase in relative exposure. Each exposure’s effect, along with its confidence interval, p-value, and the proportion of folds in which the effect was observed, are reported as follows. A1 demonstrated a significant increase in the outcome by 5.69 units, with a 95% CI of 4.81 to 6.57, identified consistently across all folds (proportion = 1.0), and a p-value < 0.01. A5 Showed a significant decrease in the outcome by -3.41 units, with a 95% CI of -4.28 to -2.54, consistently observed across all folds (proportion = 1.0), and a p-value < 0.01. A7 indicated a significant increase in the outcome by 3.40 units, with a 95% CI of 2.47 to 4.33, found in all folds (proportion = 1.0), and a p-value < 0.01. A4 had a non-significant effect of -0.36 units (inverse variance pooled), with a 95% CI of -1.69 to 0.22 and observed in 95% of folds, with p-value of 0.29. A2 exhibited a significant increase in the outcome by 1.82 units (inverse variance pooled), with 95% CIs of 0.93 to 2.72, found in 85% of folds, and both with p-values < 0.01. These findings align with the direction and relative strengths of association built into this synethetic data DGP.
We found three significant interaction effects defined as deviations when considering joint shifts in exposures compared to the sum of individual shifts. These interactions were for exposures A1-A5, A1-A7, and A2-A7, which aligns with the interactions built into this DGP. For the interaction between A1 and A5, we observe that a unit shift in A1 individually leads to an increase of 10.40 (9.08-11.73) in the outcome. On the contrary, a unit change in A5 individually shows a decrease of -2.86 (-4.10 - -1.62), suggesting an antagonistic relationship between these exposures. When both A1 and A5 are simultaneously shifted by one unit, the expected outcome under joint shift is 0.49 (-0.66 - 1.63), thus generally nullifying each exposure effect, making the interaction effect -7.26 (-8.59 - -5.92), p < 0.001. This effect was found in 45% of folds. The interaction between A1 and A7 shows that individual changes in A1 and A7 lead to increases of 14.74 (13.14 - 16.34) and 2.42 (0.62 - 4.22), respectively. However, the joint shift in X1 and X7 only increases endrocine function by 5.31 (3.65 - 6.98), resulting in an interaction effect of -11.76 (-13.4 - -10.12). This effect indicates that the joint shift is significantly less than expected under additive assumptions. This interaction was found in 20 % folds. The shift in A2 led to a 5.56 (4.19 - 6.93) and A7 a 2.83 (1.65 - 4.02), the joint shift a 4.89 (3.75 - 6.04), which led to an interaction effect of -3.86 (-5.24 - -2.48), p < 0.001, proportion of folds = 50%. Individual shifts in A5 and A7 result in changes of -2.74 (-3.70 - -1.78) and 2.61 (1.60 - 3.62), respectively. However, the interaction effect of their joint shift is -0.02 (-1.03 - 0.98), which leads to an interaction effect of -0.05 (-1.02 - 0.93), found in 70% of the folds. This suggests that these two exposures have an additive antagonism, which is built into the DGP. Due to space limitations, we do not provide fold-specific results and more details, but these are provided on github package page. Although our detection of interactions was not found in every fold (due potentially to a low sample size), we identified the majority of interactions built into these data.
6.3 Comparison to Quantile G-Computation
Quantile g-computation is a favored approach in environmental epidemiology’s mixture analyses. Estimates the joint effect of increasing all exposures by one quantile, based on linear assumptions, as: $Y_{i}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{ji}^{q}+\beta Z_{i}+\epsilon_{i}$. Here, $X^{q}$ represents the quantized components of the mixture. The methodology consists of transforming the components of the mixture into quantiles and then adding the coefficients of the linear model to produce a summary measure of the mixture, $\Psi$. However, several limitations emerge: 1. Quantiles may not accurately reflect the exposure-response relationship, which can be nonmonotonic in substances like endocrine disruptors, 2. For a comprehensible mixture estimate $\Psi$, it presumes an additive relationship of quantiles, 3. Given the intricate nature of mixtures, interactions are vital. Including interactions in quantile g-computation complicates the interpretation, as the exposure’s proportional contribution becomes variable depending on other factors’ levels, 3. This method lacks flexibility in covariate control. Our application to the NIEHS dataset with 4 quantiles, excluding interactions, yielded a scaled positive effect size of 6.28 (encompassing $A_{1},A_{2},A_{3},A_{7}$) and a negative effect size of -3.68 (involving $A_{4},A_{5},A_{6}$). Discrepancies arise when compared to the NIEHS baseline, incorrectly including $A_{3},A_{6}$. Incorporating interactions compounds the challenge. Our approach with all exposure interactions showed a joint impact of main effects of 0.02 (-3.16 - 3.20), p = 0.99 and interaction effects of 0.59 (-0.71 - 1.90), p = 0.37. Considering the parameter count and a sample size of $n=500$, it is not surprising that these estimates are not significant. However, even attempts to rigorously analyze 2-way or 3-way interactions introduce an array of complications, underscoring mixtures as an inherently data-adaptive issue and the insufficiencies of conventional methods on even straightforward synthetic datasets.

### v3

@@ 6.3 Evaluating Performance
We evaluated the asymptotic convergence to the true counterfactual differences for these exposures, in each simulation, which inherently means convergence to identifying the true exposure relationships (if bias converges to 0 for the top synergy relationship, then this means we are both finding the true synergistic relationship and correctly estimating our interaction parameter). To do so, we followed the following steps 1. From our $P_{0}$ population, we generated a random sample of size $n$ from the DGP, 2. At each iteration, we used the training sample to define the exposure(s) and create the necessary estimators for the target parameter dependent on the exposure sets. We then used the validation sample to obtain the updated causal parameter estimate using TMLE. We repeated this process for all the folds, 3. At each iteration, we output the stochastic shift estimates given the pooled TMLE for each effect, the top synergistic relationship, top antagonistic relationship, top positive marginal relationship, and top inverse marginal relationship.
To evaluate the performance of our approach, we calculated several metrics for each iteration, including bias, variance, MSE, and confidence interval (CI) coverage. To visually inspect if the rate of convergence was at least as fast as $\sqrt{n}$, we show absolute bias and MSE times $\sqrt{n}$. We then calculated the variance for each estimate and used it to calculate the mean squared error (MSE) as $\text{MSE}=\text{bias}^{2}+\text{variance}$. For each estimate, we calculated the CI coverage as the proportion of iterations the true parameter was covered by estimated confidence intervals. We calculated these performance metrics at each iteration, performing 100 iterations for each sample size $n=$ (300, 600, 1000, 2000, 3000). We used InterXshift with 5-fold cross-validation (to speed up calculations in the simulations) and default learner stacks (see below) for each nuisance parameter and data-adaptive parameter.
To use InterXshift, we need estimators for $\bar{Q}=E(Y|A,W)$ and $\xi$ (density ratio learner). InterXshift provides default algorithms to be used in a Super Learner van der Laan Mark et al. [2007] that are both fast and flexible. To estimate $\bar{Q}$, we include estimators for the Super Learner from glm, elastic net Tay et al. [2023], random forest Wright and Ziegler [2017], and xgboost Chen and Guestrin [2016]. The same learners are used for our $\xi$ learner.
@@ 6.4 Results
InterXshift consistently discovers and estimates the correct interaction. The top synergistic interaction effect was identified in 100% of the folds in all sample sizes. Figure 1 shows the absolute bias, bias times $\sqrt{n}$, CI coverage, mean squared error, and standard deviation as the sample size increases to 3000. For bias, MSE and standard deviation, there is a converge to near zero at sample size 3000. For coverage, the average coverage for each target parameter was: individual shift: 95.5%; joint shift: 96%; interaction: 96%. Coverage was slightly high at lower sample sizes but this is likely due to variability in the simulation. The absolute bias times $\sqrt{n}$ should look flat if convergence is at $\sqrt{n}$ rate which we generally see with some variability in the simulation.
Figure 1: Absolute Bias, Absolute Bias times $\sqrt{n}$, Proportion CI Coverage, Standard Deviation and MSE for Each Parameter Across Sample Sizes: These metrics represent results from CV-TMLE, that is bias is w.r.t. the oracle target parameter where we use CV to identify the estimated top ranked synergistic interaction then estimate the joint and marginal shifts for our interaction parameter.
@@ 7.1 Analysis of NIEHS Synthetic Mixtures Data
The NIEHS synthetic mixtures data is a commonly used data set to evaluate the performance of statistical methods for mixtures. These synthetic data can be considered the results of a prospective cohort study, where the outcome cannot cause exposures, and correlations between exposure variables can be thought of as caused by common sources or modes of exposure. The variable $W$ can be assumed to be a potential confounder and not a collider. The data set has 7 exposures ($A_{1}-A_{7}$) with a complex dependency structure based on endocrine disruption. Two groups of exposure ($A_{1},A_{2},A_{3}$ and $A_{5},A_{6}$) lead to high correlations within each group. $A_{1},A_{2},A_{7}$ positively contribute to the outcome, $A_{4},A_{5}$ have negative contributions, while $A_{3}$ and $A_{6}$ have no impact on the outcome. Rejecting $A_{3}$ and $A_{6}$ is difficult due to their correlations with the members of the cluster group. This correlation and effects structure is biologically plausible, as different congeners of a group of compounds may be highly correlated but have different biological effects. Exposures have various agonistic and antagonistic interactions, a breakdown of the variables sets and their relationships are: $A_{1},A_{2}$ both increase Y by concetration addition, $A_{1}-A_{4},A_{2}-A_{4},A_{1}-A_{4},A_{2}-A_{4}$ competetive antagonism, $A_{1}-A_{7},A_{2}-A_{7}$ supra-additive or synergistic, $A_{4},A_{5}$ both addivively decrease Y, $A_{4}-A_{7},A_{5}-A_{7}$ unusual antagonism. Synthetic data and the key for dataset 1 are available on GitHub. This resource shows the interactions and marginal dose relationships built into the data. Given these toxicological interactions, we expect these sets of variables to be determined in InterXshift. For example, we might expect a positive counterfactual result for A1, A2, A7 and negative results for A4 and A5. Table 1 shows the interactions built into this data.
Variables Interaction Type
A1 and A2 Toxic equivalency factor, a special case of concentration addition (both increase Y)
A1 and A4 Competitive antagonism (similarly for A2 and A4)
A1 and A5 Competitive antagonism (similarly for A2 and A4)
A1 and A7 Supra-additive (“synergy”) (similarly for A2 and A7)
A4 and A5 Toxic equivalency factor, a type of concentration addition (both decrease y)
A4 and A7 Antagonism (unusual kind) (similarly for A5 and A7)
Table 1: NIEHS Synthetic Data Interactions
Likewise, in the case for antagonistic relationships such as A1-A5 (third row in 1), we would expect a joint shift to come closer to the null, since A5 antagonizes the positive effects of A1. For A1 and A2, we would expect the joint shift to be close to the sum of individual shifts (not much interaction), but for A1 and A7 we expect to estimate more than an additive effect (some interaction). The NIEHS data set has 500 observations and 9 variables. $W$ is a binary confounder. Based on this table, we can gauge InterXshift’s performance by determining if the correct synergistic and antagonistic relationship are found and if the correct variables are rejected.
We apply InterXshift to these NIEHS synthetic data using a three-fold CV (smaller CV to save space here in k-fold result tables). We use the default libraries of learners for the estimation of $\xi$ and $Q$. We parallelize over the cross-validation to test computational run-time on a newer personal machine an analyst might be using. We use a delta of 1 for all the exposures in the mixture. We present the pooled TMLE results and the k-fold specific results.
@@ 7.2 NIEHS Synthetic Data Results
InterXshift accurately rejects exposures $A_{3},A_{6}$, finding the correct top positive marginal effects of $A_{1}$ in all the folds and the top inverse association ($A_{5}$) in all the folds. Likewise, we found the correct synergistic results built into the DGP for $A_{1}-A_{7}$ and $A_{2}-A_{7}$. In all the folds, $A_{5}-A_{7}$ was found to have the strongest antagonistic relationship, which is also true based on the DGP key provided from the NIEHS.
Table 2 presents the rank 1 positive marginal results for the NIEHS data. The first three rows display fold-specific results, with $\Psi$ representing the expected change in the outcome under the exposure shift compared to the average observed outcome. Here, $A_{1}$ is identified as the top positively associated mixture variable across all folds. The pooled TMLE column provides the aggregate results, showing an average $\Psi$ of 16.31 with a standard error of 0.37 and a 95% confidence interval of 15.59 to 17.04. These results underscore the consistent positive association of $A_{1}$ with the outcome across different data folds.
Condition $\Psi$ Variance SE Lower CI Upper CI P-value Fold N Delta
1 A1 20.66 0.45 0.67 19.35 21.98 0.00 1 167 1.00
2 A1 -2.66 1.28 1.13 -4.88 -0.44 0.02 2 167 1.00
3 A1 19.13 0.53 0.73 17.70 20.56 0.00 3 166 1.00
4 Rank 1 Pos 16.31 0.14 0.37 15.59 17.04 0.00 Pooled TMLE 500 1.00
Table 2: Rank 1 Positive Marginal Results for NIEHS Data
Condition $\Psi$ Variance SE Lower CI Upper CI P-value Fold N Delta
1 A5 -3.94 0.70 0.83 -5.57 -2.30 0.00 1 167 1.00
2 A5 -3.27 0.87 0.93 -5.10 -1.45 0.00 2 167 1.00
3 A5 -3.50 0.61 0.78 -5.04 -1.97 0.00 3 166 1.00
4 Rank 1 Neg -3.67 0.25 0.50 -4.65 -2.70 0.00 Pooled TMLE 500 1.00
Table 3: Rank 1 Negative Marginal Results for NIEHS Data
Table 3 presents the rank 1 inverse marginal results for the NIEHS data. The first three rows show fold-specific results, with $\Psi$ indicating the expected change in the outcome under the exposure shift compared to the average observed outcome. Here, $A5$ is identified as the top negatively associated mixture variable across all folds. The pooled TMLE column provides the aggregate results, demonstrating that a 1-unit increase in $A5$ leads to an average decrease of -3.67 in endocrine disruption, with a standard error of 0.50 and a 95% confidence interval of -4.65 to -2.70. These results highlight the consistent negative association of $A5$ with the outcome across different data folds.
Rank $\Psi$ Variance SE Lower CI Upper CI P-value Fold N Delta 1 Delta 2
1 1 19.94 0.44 0.66 18.64 21.25 0.00 1 167 1.00 A1
2 1 2.91 0.84 0.92 1.11 4.71 0.00 1 167 1.00 A7
3 1 25.74 2.46 1.57 21.9 26.2 0.00 1 167 1.00 A1-A7
4 1 2.89 2.03 1.43 1.39 5.30 0.00 1 167 1.00 Interaction
5 1 -2.81 1.33 1.16 -5.07 -0.54 0.01 2 167 1.00 A1
6 1 3.23 0.80 0.89 1.48 4.98 0.00 2 167 1.00 A7
7 1 24.35 0.62 0.78 22.81 25.89 0.00 2 167 1.00 A1-A7
8 1 23.93 5.03 2.24 19.53 28.32 0.00 2 167 1.00 Interaction
9 1 3.11 0.77 0.88 1.39 4.83 0.00 3 166 1.00 A2
10 1 3.72 0.44 0.66 2.42 5.02 0.00 3 166 1.00 A7
11 1 14.46 4.50 2.12 10.30 18.62 0.00 3 166 1.00 A2-A7
12 1 7.62 3.23 1.80 4.10 11.15 0.00 3 166 1.00 Interaction
Table 4: Rank 1 Synergistic Interaction Results for NIEHS Data
Table 4 presents the rank 1 synergistic interaction results for the NIEHS data. The table shows fold-specific results, with $\Psi$ indicating the expected change in the outcome under a 1-unit shift in the specified exposure(s). Here, we observe that $A1$ and $A7$ are identified as the top synergistic relationship in two folds, while $A2$ and $A7$ are identified in the remaining fold. The joint shift for $A1$ and $A7$ and for $A2$ and $A7$ reveals significant synergistic interactions built into the data-generating process. The interaction effects reflect the combined effect of shifting both exposures compared to the sum of their individual shifts, indicating strong synergy where the Psi under interaction is positive.
Rank $\Psi$ Variance SE Lower CI Upper CI P-value Fold N Delta 1 Delta 2
1 Rank 1 6.68 0.20 0.45 5.81 7.56 0.00 Pooled TMLE 500 1.00 Var 1
2 Rank 1 3.27 0.26 0.51 2.27 4.27 0.00 Pooled TMLE 500 1.00 Var 2
3 Rank 1 21.04 0.32 0.57 19.92 22.15 0.00 Pooled TMLE 500 1.00 Joint
4 Rank 1 11.08 1.01 1.00 9.11 13.05 0.00 Pooled TMLE 500 1.00 Interaction
Table 5: Pooled TMLE Estimates for Rank 1 Synergy in NIEHS Data
Table 5 presents the pooled TMLE estimates for the rank 1 synergistic relationships in the NIEHS data. The first two rows show the average individual effects of the variables involved in the top synergistic relationships, identified as Var 1 and Var 2, here because Var 1 changed across the folds it is the average across $A1$ and $A2$ and Var 2 corresponds to $A7$. The third row represents the joint effect of shifting both exposures simultaneously, while the fourth row shows the average synergistic interaction effect, which is significantly greater than the sum of individual shifts. This pooled estimate reflects the average effect across the identified top synergistic relationships.
Rank $\Psi$ Variance SE Lower CI Upper CI P-value Fold N Delta Type
1 1 -3.97 0.71 0.84 -5.62 -2.32 0.00 1 167 1.00 X5
2 1 3.00 0.84 0.91 1.21 4.79 0.00 1 167 1.00 X7
3 1 -1.76 1.00 1.00 -3.72 0.20 0.08 1 167 1.00 X5-X7
4 1 -0.78 0.93 0.96 -2.67 1.11 0.42 1 167 1.00 Interaction
5 1 -3.48 0.74 0.86 -5.16 -1.79 0.00 2 167 1.00 X5
6 1 3.21 0.80 0.89 1.46 4.96 0.00 2 167 1.00 X7
7 1 -0.61 1.22 1.10 -2.77 1.56 0.56 2 167 1.00 X5-X7
8 1 -0.34 1.08 1.04 -2.38 1.70 0.74 2 167 1.00 Interaction
9 1 -3.45 0.54 0.73 -4.89 -2.01 0.00 3 166 1.00 X5
10 1 3.83 0.43 0.66 2.54 5.13 0.00 3 166 1.00 X7
11 1 -0.69 1.17 1.08 -2.81 1.43 0.50 3 166 1.00 X5-X7
12 1 -1.07 0.90 0.95 -2.94 0.79 0.27 3 166 1.00 Interaction
Table 6: K-Fold Specific Antagonistic Rank 1 Results for NIEHS Data
Table 6 presents the k-fold specific antagonistic rank 1 results for the NIEHS data. The $\Psi$ column represents the expected change in the outcome under a 1-unit shift in the specified exposures. For the first fold, variable A5 shows a negative shift of -3.97, indicating a strong inverse association. Similarly, variable A7 shows a positive shift of 3.00. The combined shift of A5 and A7 results in an estimated change of -1.76, while the interaction term, which compares the joint shift to the two marginal shifts, shows a negative shift of -0.78. These patterns are consistent across all three folds, demonstrating significant evidence for antagonistic interactions between these variables.
7.2.1 Comparison to Existing Methods
Quantile g-computation Keil et al. [2019], prevalent in environmental epidemiology for mixture analysis, estimates the effects of uniformly increasing exposures by one quantile, based on linear model assumptions. This method quantizes mixture components, summing the linear model’s coefficients to form a summary measure ($\Psi$) for joint impact assessment. However, it inherently assumes additive, monotonic exposure-response relationships, overlooking complex, potentially nonlinear interactions typical in mixtures, like in endocrine disrupting compounds. Consequently, this method might not accurately capture the nuanced dynamics of mixed exposures, especially when interactions vary with other variable levels.
We run quantile g-computation on the NIEHS data using 4 quantiles with no interactions to investigate results using this model. The size of the scaled effect (positive direction, sum of positive coefficients) was 6.28 and included $A_{1},A_{2},A_{3},A_{7}$ and the scaled effect size (negative direction, sum of negative coefficients) was -3.68 and included $A_{4},A_{5},A_{6}$. Compared to NIEHS ground truth, $A_{3},A_{6}$ are incorrectly included in these estimates. However, the positive and negative associations for the other variables are correct. Next, because we expect interactions to exist in the mixture data, we would like to assess for them but the question is which interaction terms to include? Our best guess is to include interaction terms for all exposures. We do this and show results in Table 7.
Estimate Std. Error Lower CI Upper CI Pr($>$||||t||||)
(Intercept) 21.29 1.58 18.19 24.39 0.00
psi1 0.02 1.62 -3.16 3.20 0.99
psi2 0.59 0.67 -0.71 1.90 0.37
Table 7: Quantile G-Computation Interaction Results from NIEHS Synthetic Data
In Table 7 $\Psi_{1}$ is the summary measure for the main effects and $\Psi_{2}$ for interactions. As can be seen, when including all interactions, neither of the estimates are significant. Of course, this is to be expected given the number of parameters in the model and the sample size $n=500$. However, moving forward with interaction assessment is difficult; if we were to assess for all 2-way interaction of 7 exposures, the number of sets is 21 and with 3-way interactions is 35. We would have to run these many models and then correct for multiple testing. Hopefully, this example shows why mixtures are inherently a data-adaptive problem and why popular methods such as this, although succinct and interpretable, fall short even in a simple synthetic data set.